solve (a plane flys 2400 miles one way and takes 4 hours, then fly's back the same way and takes an hour and a half longer. what is the speed of the plane in still air and what is the wind speed?
Answer by mananth(16949)

solve (a plane flys 2400 miles one way and takes 4 hours, then fly's back the same way and takes an hour and a half longer. what is the speed of the plane in still air and what is the wind speed?
..
2400 miles = 4 hours
speed = distance /time
speed = 600 mph
..
2400 miles 5 1/2 hours = 11/2 hours
speed = distance /time
speed = 2400/(11/2)
=2400 *2/11
4800/11
=436.36 mph
..
let speed in still air be x
let speed of wind = y
...
x+y =600
x-y = 436
Add the equations
2x=1036
x= 1036/2
x= 518 mph
y = 600-518 = 82 mph